# Majority voting is not good for heaven or hell, with mirrored performance

### v4

@@ Abstract
Within the ViSE (Voting in Stochastic Environment) model, we study the effectiveness of majority voting in various environments. By the pit of losses paradox identified in previous work, majority decisions in apparently hostile environments tend to reduce the capital of society. In such cases, the simple social decision rule of “rejecting all proposals without voting” outperforms majority voting. In this paper, we identify another pit of losses appearing in favorable environments. Here, the simple social decision rule of “accepting all proposals without voting” is superior to majority voting. We prove that under a version of simple majority called symmetrized majority and the antisymmetry of the voting body, the second pit of losses is a mirror image of the pit of losses in hostile environments and explain this phenomenon. Technically, we consider a voting society consisting of individualists whose strategy is supporting all proposals that increase their capital and a group (groups) whose members vote to increase the wealth of their group. According to the main result, the expected capital gain of each agent in the environment whose generator ${X}$ has mean $\mu>0$ exceeds by $\mu$ their expected capital gain under generator $-{X}$. This result extends to location families of generators with distributions symmetric about their mean. The mentioned result determines the symmetry of the difference between the expected capital gain under the symmetrized majority and that under the “basic” social decision rule that rejects (resp. accepts) all proposals in unfavorable (resp. favorable) environments.
@@ 1.1 The problem of harmful voting
A large body of research on voting has led to the conclusion that this method of decision-making requires great care and attention when used. In particular, there are many voting paradoxes (see e.g., [36]), i.e., unexpected properties of voting, which in most cases are disappointing. A large class of paradoxes rests on a specific, carefully constructed agenda. The ViSE (Voting in Stochastic Environment) model introduces a stochastic agenda, however, this does not eliminate the paradoxical properties of majority decisions. Namely, the pit of losses paradox ([10, 26]) appears within the framework of this model. The essence of this paradox is that majority decisions in a markedly unfavorable environment are, on average, disadvantageous to society. In such cases, it is better to maintain the status quo by rejecting all proposals without a vote.
@@ 1.2 The ViSE model
The main assumptions of the ViSE model (for more details, see e.g., [10]) are as follows. A society consists of $n$ agents (also called voters). Each agent is characterized by the current real-valued capital (debt, if it is negative), which can sometimes be interpreted as utility. In each step $t=1,\ldots,T,$ some proposal is put to the vote, and the agents vote, guided by their voting strategies. In the framework of the ViSE model, a strategy is an algorithm following which an agent uses available information to decide whether to vote for, against, or abstain from voting (a ‘neutral’ vote) on the current proposal. Such a strategy may involve stochastic elements.
A proposal is a vector of algebraic capital gains of all agents. It is generated stochastically as a realization ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})={{\mathbf{{X}}}(\omega)},\,\omega\in\Omega$ of a multivariate random variable ${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ called a proposal generator. In this notation, $\omega$ is an elementary outcome, $\Omega$ is a sample space; ${\mathbf{{X}}}\!:\Omega\to{\mathbb{R}}^{n}$. We introduce $\Omega$ explicitly to distinguish it from another source of randomness arising from tie breaking.
We consider the case where the components of ${\mathbf{{X}}}$ are independent and identically distributed (i.i.d.) with a known finite mean $\mu.$ Let ${\mathcal{X}}^{\mu}$ be the set of such proposal generators. The scalar random variable ${X}{\color[rgb]{0,0,0}:\Omega\to{\mathbb{R}}}$ corresponding to a component of ${\mathbf{{X}}}$ will be called a gain generator.
The proposals put to the vote are called stochastic environment proposals. The environment is favorable if $\mu>0,$ neutral if $\mu=0,$ and unfavorable (hostile) if $\mu<0.$
@@
The dynamics of agents’ capital in various environments can be analyzed to compare voting strategies and social decision rules in order to select the optimal ones in terms of maximizing appropriate criteria. Since in this version of the model the process is stationary, it can be described by one-step characteristics. An important one is the mathematical expectation of the capital gain (expected capital gain), ECG, of an agent with a certain strategy after implementing one social decision of a given society in a given environment. The ECG can be negative, zero, or positive.
@@ 1.3 Relation between the ViSE model and other models
Proposals generated exogenously and having random effects were considered by [38] and [13, 14]111[14] emphasized the importance of analyzing dynamic voting in stochastic environments and noted the scarcity of relevant work. The ViSE model ([8], etc.) differs significantly from the Dziuda and Loeper model, but belongs to the same class., but the non-transferable nature of the individual utilities (i.e., measuring utilities in subjective scales rather than in a common ‘currency’) made it difficult to study cooperative and prosocial strategies within the models used. A simulation study of majority voting by bots with a random agenda in a model with ideal points was conducted by [6].
Research on cooperation under dynamic voting in the pie-sharing game with transferable utilities was initiated by [16]. [15] performed an experimental study of cooperative solutions. [20] obtained interesting results on coalition dynamics and decision effectiveness in a model with discounting and bribes. The results of [10], where we looked for the optimal threshold of qualified majority, are comparable with the results of [24], where a stochastic model was studied for which, under certain conditions, simple majority was optimal. The problem of the optimal voting threshold was also considered by [11], where proposals were generated exogenously and the profitability of the first accepted proposal was evaluated; in that model, the unanimity rule would be optimal if there were no discounting.
@@
The analysis of effectiveness brings this framework closer to the one studied by [22], where there is no dynamics, since a number of (exogenously generated) proposals are voted ‘in parallel’, but the acceptance of any proposal yields some change in individual utilities, and the characteristics of the resulting gain vector allow one to evaluate the effectiveness of the decision-making mechanism.
@@
It is worth noting that decision making within the ViSE model can be formalized as a general Markov decision process (see, e.g., [18, 41, 17]), where states are vectors of the current capital of the agents, and the possible social decision rules are as follows: vote by simple majority, vote using some other rules, accept the submitted proposal without voting, and reject the proposal without a vote. The last two rules can be treated as degenerate voting rules, namely, the threshold majorities with a negative threshold and with a threshold exceeding the number of voters, respectively. The probability (or density) of the transition between states $s$ and $s^{\prime}$ is determined by the social decision rule and the probability (density) associated by the proposal generator with the proposal $s^{\prime}-s$. The society’s reward is the sum of the components of an accepted proposal, or zero if the proposal is rejected. This perspective is useful for interpreting results obtained within the ViSE model (cf. Remark 1).
1.4 Pit of losses paradox
@@
The ViSE model leaves no room for a purposeful agenda: it is replaced by a favorable or unfavorable stochastic environment that generates proposals. However, a voting paradox of the above nature still arises. Namely, by [10, Proposition 1], if $\mu,$ the mean of a Gaussian gain generator ${X}$, is below a certain negative threshold (which is proportional to the standard deviation of ${X}$ and depends on the number of voters), then for a purely individualistic society and majority rule, the expected reward (one-step capital gain, ECG) of each agent is also negative. This is the pit of losses paradox illustrated in Fig. 1a. It means that sequential voting will inevitably bankrupt all participants in such a society.
@@
Figure 1: The expected capital gain (ECG) of an agent in a voting society of 21 individualists: (a) in unfavorable environments $(\mu<0)$; (b) in some favorable environments $(\mu>0)$
This implies that in some unfavorable environments, majority rule is worse for each agent than rejecting all proposals without a vote, which saves capital. For some gain generators distributed with longer tails than Gaussian generators, the pit of losses also exists, but it is shallower.
Thus, the essence of the paradox contained in the chaos theorems [^2] It should be noted, however, that sometimes decision making routines that encourage chaotic conflict are considered to be robust to bounded rationality and “promote effective search in an uncertain task environment” ([19]). Furthermore, chaos is limited because with a random agenda, some areas are more attractive than others (see, e.g., [6]). is not agenda manipulation, but the combination of selfishness with majority rule. Methods to resolve the pit of losses paradox include cooperation ([9]), prosocial behavior ([43, 44]) that can be supported by income redistribution through taxes ([1]), and optimizing the voting threshold ([10, 26]; see Section 4 below).
@@ 1.5 Contributions of this study
The pit of losses paradox is characteristic of unfavorable environments (Fig. 1a). What happens in favorable environments? One might assume (cf. Fig. 1b) that the agent’s expected capital gain curve approaches the line $\mathrm{ECG}=\mu$ in the first quadrant from above as $\mu$ increases, i.e., that majority decisions provide some decreasing positive gain over what the environment offers.
However, somewhat surprisingly, this curve intersects the line $\mathrm{ECG}=\mu$ (or $\mathrm{ECG}/\sigma=\mu/\sigma$, as shown in Fig. 1b). This means that under sufficiently favorable conditions, filtering proposals by majority vote is harmful. Furthermore, the abscissa of the intersection is equal in absolute value to the zero of the same function, marked with a black circle in Fig. 1a. In Sections 3 and 4 we study and explain this phenomenon in the more general case of a society consisting of individualists and a group (or groups).
Let us consider the following agent strategies. An individualist casts a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) their own capital. The ‘half a vote’ is a reasonable contribution in case of abstention. It ensures the complementarity effect of opposite proposals for discrete/discontinuous distributions in the results below. The voting strategy of a group member is to cast a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) the total capital of the group this voter belongs to. Such a group can also be called a utilitarian group. Clearly, an individualist is a special case of a group.
@@
Lemma 1 below shows that under some conditions and any generator ${\mathbf{{X}}}$ with mean $\mu>0$, the expected capital gain of any agent exceeds that under generator $-{\mathbf{{X}}}$ by $\mu.$ In Corollary 6, this result is applied to voting bodies consisting of individualists and group members. Corollary 8, an easy consequence of the previous results, states that the same difference of $\mu$ holds for generators ${\mathbf{{X}}}^{\mu}$ with mean $\mu$ and ${\mathbf{{X}}}^{-\mu}={\mathbf{{X}}}^{\mu}-2\mu,$ provided that ${\mathbf{{X}}}^{\mu}$ is symmetric about $\mu.$ In Section 4 we introduce the concept of expected relative capital gain (ERCG). This value is the difference between the ECG under majority rule and that under the ‘basic’ social decision rule that rejects (resp. accepts) all proposals in unfavorable (resp. favorable) environments. We show that under the conditions of Corollary 8, the ERCG of any agent is an even function of $\mu,$ i.e., $\mbox{ERCG}(-\mu)=\mbox{ERCG}(\mu)$ for any $\mu$. This allows us to conclude that the pit of loses paradox extends to favorable environments, and the performance of the symmetrized majority rule is symmetric about 00. That is why “majority voting is not good for heaven or hell, with mirrored performance” (see Section 4).
In the above paragraph, we presented Corollaries 6 and 8, while Theorems 5 and 7 hold for the more general class of voting bodies formed by agents with ‘complementary’ strategies and a ‘complementary’ voting rule.
@@ 2 Basic notation and problem formulation
We consider a society in which $\ell$ agents are individualists, while $g$ agents form a group, so that $g+\ell=n$ is the size of the society. Let us say that the set of agents with their voting strategies and an adopted social decision rule form a voting body.
@@
Given a voting body, let $n^{+}({\mathbf{{x}}})$ be the number of votes cast for proposal ${\mathbf{{x}}}$; let $I({\mathbf{{x}}})$ be the probability of accepting (and thus implementing) ${\mathbf{{x}}}$. If the social decision rule and the agents’ strategies are not stochastic, then, for any ${\mathbf{{x}}},$ $I({\mathbf{{x}}})$ is either 00 or $1$
We say that a voting body is antisymmetric (or satisfies reversal symmetry; cf. [7]) if for any proposal ${\mathbf{{x}}},$ the probabilities of ${\mathbf{{x}}}$ and $-{\mathbf{{x}}}$ being accepted sum to $1$:
[…]
@@
Let ${\mathbf{D}}_{\mathbf{{X}}}$ denote the random vector of capital gains that implements the social decisions on the proposals generated by ${\mathbf{{X}}}.$ We note that ${\mathbf{D}}_{\mathbf{{X}}}$ is random due to its dependence on both ${\mathbf{{X}}}$ and the tie-breaking outcome. Therefore, the sample space of ${\mathbf{D}}_{\mathbf{{X}}}$ is the Cartesian product of $\Omega$ and the sample space of the tie-breaking rule.
@@
$\{1,\ldots,g\}$ group; $0\leq g\leq n$
$\{g+1,\ldots,n\}$ set of individualists; $\ell=n-g$ is the number of them
@@
$\mu=\mu\mathop{\!{}_{{X}}}={\mathbb{E}}({X})$ mean of ${X}$
$\sigma=\sigma\mathop{\!{}_{{X}}}$ standard deviation of ${X}$
@@
${\mathcal{X}}^{\mu}$ set of proposal generators whose components are i.i.d.
gain generators with mean $\mu$
${\mathbf{{X}}}^{\mu}{\color[rgb]{0,0,0}\,\in{\mathcal{X}}^{\mu}}$ proposal generator corresponding to gain generator ${X}^{\mu}$
@@
The purpose of this study is to find out how the effectiveness of majority decisions depends on the favorability of the environment. We examine the relationship between ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{1}}\bigr{)}$ and ${\mathbb{E}}\bigl{(}{\mathbf{D}}_{{\mathbf{{X}}}_{2}}\bigr{)}$ for opposite proposal generators ${\mathbf{{X}}}_{1}$ and ${\mathbf{{X}}}_{2}$, with two different notions of opposition. Numerical examples correspond to the normal distribution.
3 Main results

### v5

@@ Abstract
Within the ViSE (Voting in Stochastic Environment) model, we study the effectiveness of majority voting in various environments. As shown by the pit-of-losses paradox identified in previous work, majority decisions in apparently hostile environments tend to reduce the capital of society. In such cases, the simple social decision rule of “rejecting all proposals without voting” outperforms majority voting. In this paper, we identify another pit of losses appearing in favorable environments; here, the simple social decision rule of “accepting all proposals without voting” is superior to majority voting. We prove that, under a version of simple majority called symmetrized majority and under the antisymmetry of the voting body, this second pit of losses is a mirror image of the one arising in hostile environments, and we explain this phenomenon. Technically, we consider a voting society consisting of individualists who support all proposals that increase their personal capital and a group (or groups) whose members vote to increase their group’s wealth. According to the key lemma, the expected capital gain of each agent under the social decision rule when the random gain generator is X{X} with mean $\mu>0$ exceeds their expected gain under the reflected generator X-{X} by exactly $\mu$. This extends to location-scale families of generators with distributions symmetric about their mean. This result reveals a mirror symmetry in the performance of the symmetrized majority rule relative to a baseline rule. The baseline rule accepts all proposals in favorable environments and rejects them in unfavorable (hostile) ones.
@@ 1.1 The problem of harmful voting
A large body of research on voting has led to the conclusion that this method of decision-making requires great care and attention when used. In particular, there are many voting paradoxes (see e.g., [36]), i.e., unexpected properties of voting, which in most cases are disappointing. A large class of paradoxes rests on a specific, carefully constructed agenda. The ViSE (Voting in Stochastic Environment) model introduces a stochastic agenda, however, this does not eliminate the paradoxical properties of majority decisions. Namely, the pit-of-losses paradox ([10, 26]) appears within the framework of this model. The essence of this paradox is that majority decisions in a markedly unfavorable environment are, on average, disadvantageous to society. In such cases, it is better to maintain the status quo by rejecting all proposals without a vote.
@@ 1.2 The ViSE model
The main assumptions of the ViSE model (for more details, see e.g., [10]) are as follows. A society consists of $n$ agents (also called voters). Each agent is characterized by the current real-valued capital (debt, if it is negative), which can sometimes be interpreted as utility. In each step $t=1,\ldots,T,$ some proposal is put to the vote, and the agents vote, guided by their voting strategies. In the framework of the ViSE model, a strategy is an algorithm following which an agent uses available information to decide whether to vote for, against, or abstain from voting (a “neutral” vote) on the current proposal. Such a strategy may involve stochastic elements.
A proposal is a vector of algebraic capital gains of all agents. It is generated stochastically as a realization ${\mathbf{{x}}}=({{x}_{1}},\ldots,{{x}_{n}})$ of a multivariate random vector ${\mathbf{{X}}}=({{X}_{1}},\ldots,{{X}_{n}})$ called a proposal generator.
We restrict to the case where the components of $\mathbf{{X}}$} are independent and identically distributed (i.i.d.) with a known finite mean $\mu.$ Let ${{\mathcal{X}}}^{\mu}$ be the set of such proposal generators. A scalar random variable X{X} that denotes an arbitrary component ${X}_{i}$} of $\mathbf{{X}}$} will be called a gain generator.
In the ViSE model, the proposals put to the vote are also called stochastic environment proposals. The environment is favorable if $\mu>0,$ neutral if $\mu=0,$ and unfavorable (or hostile, in intuitive terms) if $\mu<0.$
@@
The dynamics of agents’ capital in various environments can be analyzed to compare voting strategies and social decision rules in order to select the optimal ones in terms of maximizing appropriate criteria. In this paper, we consider a variant of the model where the process is stationary, that is, it can be described by single-step characteristics. An important one is the mathematical expectation of the capital gain (expected capital gain), ECG, of an agent with a certain strategy after implementing one social decision of a given society in a given environment. The ECG can be negative, zero, or positive. On the other hand, we view this decision as a step in various multi-step processes that are not necessarily stationary due to their additional features.
@@ 1.3 Relation between the ViSE model and other models
Proposals generated exogenously and having random effects were considered by [38] and [13, 14]111[14] emphasized the importance of analyzing dynamic voting in stochastic environments and noted the scarcity of relevant work. The ViSE model ([8], etc.) differs significantly from the Dziuda and Loeper model, but belongs to the same class., but the non-transferable nature of the individual utilities (i.e., measuring utilities in subjective scales rather than in a common “currency”) made it difficult to study cooperative and prosocial strategies within the models used. A simulation study of majority voting by bots with a random agenda in a model with ideal points was conducted by [6].
Research on cooperation under dynamic voting in the pie-sharing game with transferable utilities was initiated by [16]. [15] performed an experimental study of cooperative solutions. [20] obtained interesting results on coalition dynamics and decision effectiveness in a model with discounting and bribes. The results of [10], where we looked for the optimal threshold rule, are comparable with the results of [24], where a stochastic model was studied for which, under certain conditions, simple majority was optimal. The problem of the optimal voting threshold was also considered by [11], where proposals were generated exogenously and the profitability of the first accepted proposal was evaluated; in that model, the unanimity rule would be optimal if there were no discounting.
@@
The analysis of effectiveness brings this framework closer to the one studied by [22], where there is no dynamics, since a number of (exogenously generated) proposals are voted “in parallel”, but the acceptance of any proposal yields some change in individual utilities, and the characteristics of the resulting gain vector allow one to evaluate the effectiveness of the decision-making mechanism.
@@
It is worth noting that decision making within the ViSE model can be formalized as a general Markov decision process (see, e.g., [18, 41, 17]), where states are vectors of the current capital of the agents, and the possible social decision rules are as follows: vote by simple majority, vote using some other rules, accept the submitted proposal without voting, and reject the proposal without a vote. The last two rules can be treated as degenerate voting rules, namely, the threshold rules with a negative threshold and with a threshold exceeding the number of voters, respectively. The probability (or density) of the transition between states $s$ and $s^{\prime}$ is determined by the social decision rule and the probability (density) associated by the proposal generator with the proposal $s^{\prime}-s$. The society’s reward is the sum of the components of an accepted proposal, or zero if the proposal is rejected. This perspective is useful for interpreting results obtained within the ViSE model (cf. Remark 1).
1.4 Pit-of-losses paradox
@@
The ViSE model leaves no room for a purposeful agenda: it is replaced by a favorable or hostile stochastic environment that generates proposals. However, a voting paradox of the above nature still arises. Namely, by [10, Proposition 1], if $\mu,$ the mean of a Gaussian gain generator X{X}, is below a certain negative threshold (which is proportional to the standard deviation of X{X} and depends on the number of voters), then for a purely individualistic society and majority rule, the expected reward (one-step capital gain, ECG) of each agent is also negative. This is the pit-of-losses paradox illustrated in Fig. 1a. It means that sequential voting will inevitably bankrupt all participants in such a society.
@@
Figure 1: The expected capital gain (ECG) of an agent in a voting society of 21 individualists: (a) in unfavorable (hostile) environments $(\mu<0)$; (b) in some favorable environments $(\mu>0)$
This implies that in some hostile environments, majority rule is worse for each agent than rejecting all proposals without a vote, which saves capital. For some gain generators distributed with longer tails than Gaussian generators, the pit of losses also exists, but it is shallower.
Thus, the essence of the paradox contained in the chaos theorems [^2] It should be noted, however, that sometimes decision making routines that encourage chaotic conflict are considered to be robust to bounded rationality and “promote effective search in an uncertain task environment” ([19]). Furthermore, chaos is limited because with a random agenda, some areas are more attractive than others (see, e.g., [6]). is not agenda manipulation, but the combination of selfishness with majority rule. Methods to resolve the pit-of-losses paradox include cooperation ([9]), prosocial behavior ([43, 44]) that can be supported by income redistribution through taxes ([1]), and optimizing the voting threshold ([10, 26]; see Section 4 below).
@@ 1.5 Contributions of this study
The pit-of-losses paradox is characteristic of hostile environments (Fig. 1a). What happens in favorable environments? One might assume (cf. Fig. 1b) that the agent’s expected capital gain curve approaches the line $\mathrm{ECG}=\mu$ in the first quadrant from above as $\mu$ increases, i.e., that majority decisions provide some decreasing positive gain over what the environment offers.
However, somewhat surprisingly, this curve intersects the line $\mathrm{ECG}=\mu$ (or $\mathrm{ECG}/\sigma=\mu/\sigma$, as shown in Fig. 1b). This means that under sufficiently favorable conditions, filtering proposals by majority vote is harmful. Furthermore, the abscissa of the intersection is equal in absolute value to the zero of the same function, marked with a black circle in Fig. 1a.
Let us consider the following agent strategies. An individualist casts a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) their own capital. The “half a vote” is a reasonable contribution in case of abstention. It ensures the complementarity effect of opposite proposals for discrete/discontinuous distributions in the results below. The voting strategy of a group member is to cast a vote (half a vote) for a proposal whenever this proposal increases (respectively, preserves) the total capital of the group this voter belongs to. Such a group can also be called a utilitarian group. Clearly, an individualist is a special case of a group.
@@
Lemma 1 below shows that, under certain conditions and for a generator $\mathbf{{X}}$} with mean $\mu>0$, the expected capital gain of any agent exceeds that under the generator $\mathbf{{X}}$} exactly by $\mu.$ In Corollary 7, this result is applied to voting bodies formed by individualists, group members, and the symmetrized majority rule. Corollary 9 states that, for any location-scale family of gain generators ${X}^{\mu}={X}^{0}+\,\mu,\,\mu\in{\mathbb{R}},$ with symmetric ${X}^{0},$ the same difference of $\mu$ holds for the elements ${\mathbf{{X}}}^{\mu}$ and ${\mathbf{{X}}}^{-\hskip-0.90138pt\mu}$.
Corollaries 7 and 9 are special cases of Theorem 6 and Corollary 8, which apply to voting bodies formed by agents with complementary strategies and a complementary social decision rule and to all antisymmetric voting bodies, respectively. According to Corollary 10, for location-scale families of gain generators with symmetric distributions and antisymmetric voting bodies, the performance of a social decision rule relative to the baseline rule is an even function of $\mu.$ The baseline rule accepts all proposals in favorable environments and rejects them in unfavorable ones.
Thus, in the case of complementary voter strategies, the performance of the symmetrized majority rule is symmetric about $\mu\!=\!0$. Thereby, the pit of loses paradox extends to favorable environments; in other words, “majority voting is not good for heaven or hell, with mirrored performance”. Finally, we consider the optimal threshold rule and show that its performance in individualistic societies also exhibits mirror symmetry.
@@ 2 Basic notation and problem formulation
We consider a society in which some agents are individualists, while other agents form one or several groups. Let us say that the set of agents with their voting strategies and an adopted social decision rule form a voting body.
@@
Suppose that the only information that agents’ strategies use is the current proposal $\mathbf{{x}}$}.
Given a voting body, let $n^{+}({\mathbf{{x}}})$ be the number of votes cast for proposal $\mathbf{{x}}$}; let $I({\mathbf{{x}}},{\mathbf{{X}}})$ be the probability of accepting (and thus implementing) $\mathbf{{x}}$} under $\mathbf{{X}}$} by the voting body. If the social decision rule and the agents’ strategies are not stochastic, then, for any $\mathbf{{x}}$} and $\mathbf{{X}}$}, $I({\mathbf{{x}}},{\mathbf{{X}}})$ is either 0 or $1.$
We say that a voting body is antisymmetric (or satisfies reversal symmetry; cf. [7]) if for any generator $\mathbf{{X}}$} and proposal $\mathbf{{x}}$} it generated, the probabilities of $\mathbf{{x}}$} and $\mathbf{{x}}$} being accepted under $\mathbf{{X}}$} and $\mathbf{{X}}$}, respectively, sum to $1$:
$\color[rgb]{0,0,0}I({\mathbf{{x}}},{\mathbf{{X}}})+I(-{\mathbf{{x}}},-{\mathbf{{X}}})=1\mbox{~~holds for all~~}{\mathbf{{x}}}\in{\mathbb{R}}^{n}\mbox{~~and all~~}{\mathbf{{X}}}\in{\mathcal{X}}.$} (1)
@@
Let ${\mathbf{D}}_{\mathbf{{X}}}$ denote the random vector of capital gains that implements the social decisions on the proposals generated by $\mathbf{{X}}$} (${\mathbf{D}}_{\mathbf{{X}}}=\bm{0}$ when ${\mathbf{{X}}}={\mathbf{{x}}}$ is rejected). We note that ${\mathbf{D}}_{\mathbf{{X}}}$ is random due to its dependence on both $\mathbf{{X}}$} and the tie-breaking outcome.
@@
$\mu=\mu\mathop{\!{}_{{X}}}={\mathbb{E}}({X})$ mean of X{X}
$\sigma=\sigma\mathop{\!{}_{\hskip-0.90138pt{X}}}$ standard deviation of X{X}
@@
$\mathcal{X}$} set of proposal generators whose components are i.i.d.
${\mathbf{{X}}}^{\mu}$ proposal generator corresponding to gain generator ${X}^{\mu}$
${{\mathcal{X}}}^{\mu}$ set of proposal generators with mean component value $\mu$
@@
The purpose of this study is to find out how the effectiveness of social decisions depends on the favorability of the environment. We examine the relationship between ${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}_{1}}\bigr)$ and ${\mathbb{E}}\bigl({\mathbf{D}}_{{\mathbf{{X}}}_{2}}\bigr)$ for opposite proposal generators ${\mathbf{{X}}}_{1}$ and ${\mathbf{{X}}}_{2}$, with two different notions of opposition. Numerical examples correspond to the normal distribution.
3 Antisymmetric voting bodies